OpenStudy (anonymous):

What is the quotient written in scientific notation? 2x10^21 -------- 8x10^3 A)2.5x10^17 B)2.5x10^6 C)0.25x10^18 D)0.25x10^7

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no i'm bad at this

OpenStudy (skullpatrol):

The answer is C, do you want to know why?

OpenStudy (anonymous):

yes please and thank you (:

OpenStudy (skullpatrol):

The property of quotients says (ac)/(bd) = (a/b)•(c/d) Do you see how that^ formula applies to this question?

OpenStudy (skullpatrol):

$$\Huge\dfrac{ac}{bd}=\dfrac{a}{b}•\dfrac{c}{d}$$

OpenStudy (anonymous):

ohh yeah thanks

OpenStudy (skullpatrol):

Thanks for trying to learn and understand :)

OpenStudy (anonymous):

thank you so much for your help and teaching me (:

OpenStudy (skullpatrol):

The property of quotients says (ac)/(bd) = (a/b)•(c/d) Do you see how that^ formula applies to this question?